# Efficient Cloud-Edge Collaborative Inference for Object Re-Identification

### v1

@@ 4.3 Objective Function
In this part, we describe how to train the DaCM network. For the training set $\mathcal{D}^{\mathrm{train}}$, we randomly select two samples ($q$ and $p$) with the same label ID but different camera IDs, and their spatial-temporal information is denoted $\{l^{q},t^{q},c^{q}\}$ and $\{l^{p},t^{p},c^{p}\}$, respectively. According to the time difference $t^{p}-t^{q}$ and camera id $c^{q}$, we obtain the output $\mathbf{y}$ via the proposed network, and then we optimize the network via the expectation:
$\arg\max_{\theta}\mathbb{E}_{(q,p)\sim\mathcal{D}^{\mathrm{train}}}[P(y=c^{p}|t^{q},t^{p},c^{q};\theta)],$ (7)
where $\theta$ is the parameters of the network. In implementation, this expectation can be easily converted to a classification problem, and we adopt the cross-entropy loss function to generate the backward gradients:
$\ell=-\log\frac{\exp(\mathbf{y}_{c^{p}})}{\sum_{i}^{C}\exp(\mathbf{y}_{i})}.$ (8)
Actually, the classification solves such a problem: determining the camera at which an object will appear after a duration of $t^{p}-t^{q}$, starting from camera $c^{q}$. the network will acquire knowledge about the topology of devices deployed in the system and some characteristics of the target’s movement, thereby facilitating efficient inference.
@@ 4.4 Inference
As discussed in Sec. 3, to achieve efficient inference, the system should promptly return the desired image to the cloud server by generating appropriate $e$ and $b$. In this section, we elaborate on how to accomplish this goal using the DaCM network.
@@
$b^{i}=\frac{z^{i}}{\sum_{i}z^{i}}*B,~{}z^{i}=\phi\left(\frac{\mathbf{y}}{\gamma_{0}}\right)_{i}*\left(\frac{\exp(|\mathcal{G}^{i}|)}{\gamma_{1}\sum_{j}\exp(|\mathcal{G}^{j}|)}\right),$ (10)
where $\phi$ is the $\operatorname{softmax}$ function. Note that for cloud-level inference, $t^{d}$ should be provided by the user. If there is no such information, and DaCM network does not operate in the cloud.
Edge-level inference. For cloud-level inference, the focus is on re-ranking the index of gallery images in the uploading sequence by generating proper $s_{j}^{i}$ as discussed in Eq. (1). Denoted the query image with its spatial-temporal information as $\{\mathcal{I}^{q},c^{q},t^{q}\}$, one gallery image at the $c^{d}$-th edge device as $\{\mathcal{I}^{d},c^{d},t^{d}\}$, we send $\{c^{q},t^{q},t^{d}\}$ into DaCM and obtain the output $\mathbf{y}$. If $\mathbf{y}_{c^{q}}$ is small, it implies that $\mathcal{I}^{q}$ and $\mathcal{I}^{d}$ do not match in spatio-temporal correlation, resulting in a minimal likelihood of having the same ID as the query image. This process can be simpley formulated as $\operatorname{softmax}(\mathbf{y})_{c^{d}}$. Applying this operation to all gallery images on the $i$-th edge device, we obtain spatial-temporal similarity. Next, we delve into how to combine it with visual similarity.
Given a reliable visual similarity, it is difficult to build a reliable joint metric because the spatial-temporal similarity is unreliable and it is hard to assign appropriate weighting factors for these two types of metrics. Inspired by the joint metric proposed in [33], we adopt a smoothing operator to alleviate unreliable probability estimation. Denoted the spatial-temporal similarity as $\mathbf{o}\in\mathbb{R}^{N_{i}}$ and the visual similarity as $\mathbf{v}\in\mathbb{R}^{N_{i}}$, the joint similarity can be computed as:
$s_{k}^{i}$ $=\eta\left(\mathbf{\hat{o}}_{k};\alpha\right)\eta\left(\mathbf{v}_{k};\alpha\right),$ (11)
$\eta(x;\alpha,\beta)$ $=\frac{1}{1+\exp(-\alpha x)},$
$\mathbf{\hat{x}}$ $=\operatorname{softmax}(\mathbf{x}/\beta).$
Finally, the gallery images are re-ranked according to $\mathbf{s}^{i}$.
Time-constrained ReID For a ReID system, users sometimes wish to search for targets near a specific time, a task challenging to accomplish solely based on visual features. Sorting only by time may introduce a large number of unrelated images. Therefore, the key to achieving tcReID lies in how to effectively combine time information with visual information. We observe that Eq. (11) offers a natural way to fulfill such a task. However, Eq. (11) does not satisfy the requirement because it does not introduce $t^{d}$ into the similarity calculation. Therefore, we propose a new formulation to meet the requirements of tcREID task.
@@
$\hat{s}_{k}^{i}=s_{k}^{i}\eta\left(\cos(\mathbf{B}_{k},\mathbf{y})\right),$ (12)
where $\cos$ denotes the cosine distance function. Finally, we sort the gallery images according to $\hat{\mathbf{s}}^{i}$ and return them to the cloud server in batches.
@@ 5.1 Experimental Settings
Datasets. We mainly evaluate our proposed framework and method on the DukeMTMC-reID [39] dataset, since it is annotated with high-quality timestamp. It is a subset of DukeMTMC [27] for Person ReID, which has 1,404 identities across multiple cameras, with 408 ”distractor” identities. It’s split into 702 training and 702 testing IDs. The training set has 16,522 images, and the testing set includes 2,228 query images and 17,661 gallery images.
Compared methods: We first compare the proposed approach with several inference strategies to show its performance, including:
  • Centralized denotes the conventional centralized inference strategy, which collects all images captured by edge devices and conducts similarities calculations in the cloud server. We use it as the baseline to show the boosting effectiveness of different inference strategies.
  • Baseline denotes the simple cloud-edge collaborative inference strategy: the amount of transmission is evenly distributed to each edge device and each edge device assigns the upload sequence according to the distance between the query image and gallery images. By comparing with this strategy, we can see the improvements brought by the proposed DaCM network.
  • stReID [33] interpolates the statistical spatial-temporal distribution into the similarity calculation process for person ReID. Thus, we can introduce it to the cloud-edge collaborative framework and treat it as one kind of edge-level inference strategies to re-rank the gallery images.
  • Sequence and Pairwise are two different strategies to compute the spatial-temporal distribution, which are similar to stReID, and we also adopt them as the edge-level inference strategies.
  • Spatula [16] uses statistical spatial-temporal distributions to prune the video frames for object ReID. Thus, we can adopt it as one kind of cloud-level inference strategies to adjust the amount of transmission of each edge device.
  • Ours(OC), Ours(OE), and Ours are all our proposed methods. OC and OE denote only cloud-level utilization and only edge-level utilization, respectively. Such comparison can show the effect of these two kinds of utilization applied by the proposed DaCM network.
More details of these strategies can be found in the supplementary materials.
Hyper-parameters: We employ Adam as the optimizer and set its weight decay to 1e-5. The initial learning rate is set to 0.01 and is reduced by 10 for every 30 epochs. $\gamma_{0}$ and $\gamma_{1}$ are both set to 0.01 as the default. $\alpha$ and $\beta$ are bot set to 0.1. $\lambda$ in Eq. (3) is set to 10,000 as the default. $B$ is set to $3*C$, i.e., each edge device can upload an average of three images at a time. More analysis of such hyper-parameters please refer to our supplementary materials.
@@ 5.2 Evaluation Protocols
In this part, we first introduce two evaluation protocols (R-K and mAP) that are widely used in previous ReID methods, and then we propose several novel protocols (pR-K, mpR, mTN, and LoA) to show the efficiency and effectiveness of the cloud-edge collaborative inference.
Conventional protocols:  For each query, one distance function is used to compute the similarities between the query image and all the gallery images and return a ranked table from small to large. For each individual query identity, his/her/its gallery samples from the same camera are excluded due to the setting of cross-view matching in ReID. Rank@K (R-K) is obtained by checking if top-k gallery images contain the query identity. Mean average precision (mAP) is used to evaluate the overall performance. For each query, we calculate the area under the Precision-Recall curve, i.e., average precision (AP). Then, the mean value of APs of all queries, i.e., mAP, is calculated, which considers both precision and recall of a ReID method, thus providing a more comprehensive evaluation.
Proposed protocols:  Let us initially provide a definition of the desired image, as the proposed protocols hinge upon this conceptual foundation. A desired image is a particular sample among the gallery images, sharing the same identity as the query image and possessing a timestamp in proximity to a given target time.
  • mean Transmission Number(mTN) is a protocol used to present the efficiency of the method. For each query image, there exists one corresponding TN, and we average the TN of all query images as mTN.
  • precise Rank@K (pR-K) is calculated by checking whether top-k gallery images contain the desired image that has the same ID with the query image and is closest to the target time, so pR-K is a stricter protocol than R-K.
  • mean precise Rank (mpR) For the $i$-th query data, when pR-$k_{i}$ is successful but pR-($k_{i}$-1) is not successful, we record its precise Rank as $k_{i}$, and we average the $k_{i}$ of all query data as mpR.
In summary, R-K and mAp are used to evaluate the performance of the normal ReID task, mTN is used to evaluate the efficiency of the ReID system, and pR-K and mpR are used to evaluate the performance of the tcReID task. For more details of protocols refer to the supplementary materials.
5.3 Experimential Results
@@
Centralized 79.86 90.75 96.77 1561 0.94 123.77
@@
[…]
Ours(OE) 89.84 96.68 98.20 4.43 44.74 1.48
Ours 89.84 96.68 98.20 3.25 44.74 1.48
Ours(Lin) 62.43 69.21 73.92 7.05 20.38 20.56
Ours(Sam) 74.16 81.73 86.57 4.76 23.37 9.35
Table 2: Boosting Effect evaluated on DukeMTMC-reID dataset.
[…]
BoW+kissme [38] 25.1 - - 12.2
LOMO+XQDA [20] 30.8 - - 17.0
PAN [40] 71.6 - - 51.5
SVDNet [29] 76.7 - - 56.8
HA-CNN 80.5 - - 63.8
APR [21] 70.7 - - 51.9
Human Parsing [18] 84.4 91.9 93.7 71.0
PSE+ECN [28] 85.2 - - 79.8
MultiScale [3] 79.2 - - 60.6
CLIP-ReID [19] 90.0 - - 80.7
PCB [30] 83.8 91.7 94.4 69.4
PCB + st-ReID [33] 94.0 97.0 97.8 82.8
PCB + InSTD [26] 92.7 96.5 97.6 86.1
PCB + Ours(OE) 96.2 98.3 99.0 89.5
SBS [10] 90.8 95.5 96.8 79.9
SBS + st-ReID 93.7 96.9 97.7 82.2
SBS + Ours(OE) 96.7 98.2 99.0 89.8
TranReID [11] 90.8 95.7 97.3 81.8
TranReID + st-ReID 92.4 96.4 97.5 83.4
TranReID + Ours(OE) 96.8 98.3 98.7 91.0
Efficiency of the proposed ReID system. Firstly, we compare the proposed approach with the strategies introduced in Sec. 5.1, and the results are shown in Table 1. By analyzing the mTN values of different strategies, we can see that Centralized obtains a huge number of mTN since it requires uploading all images to the cloud server, and a simple cloud-edge collaborative framework (Baseline) reduces mTN to 9.56, which saves much network traffic. stReID, Sequence, and Pairwise also yield more promising results than the baseline, but our method achieves the lowest mTN among these methods, demonstrating its efficiency. Meanwhile, the results in the table also show that using DaCM network alone in the cloud server or in the edge devices can reduce the mTN to a certain extent (from 9.56 to 5.39 and 4.43, respectively), and the combination of them can lead to an optimal result. In addition, we also present the results produced by different spatial-temporal embedding methods (Linear, Sampling, and our used sinusoidal). We can see that the difference in the embedding method can have a huge impact on the results and the sinusoidal embedding used gives the best results.
As for the conventional protocols (mAP/R-1), we can see that stReID boosts the performance from 79.86/90.75 to 82.19/96.95, which is a huge improvement. Our method can further improve the performance of mAP and R-K (89.84/96.58), demonstrating that learning to model the spatial-temporal correlation is better than just simple statistics. As for the protocols of pR-1 and mpR, they are designed to evaluate the performance of the tcReID task, but most of the previous methods do not take it into consideration, and their methods only produce meaningless output. For example, if we use the visual similarity for tcReID, it only achieves 0.94 pR-1. However, we can learn that our method can achieve 44.74 pR-1, which is still low but it demonstrates that the proposed approach can be applied to the challenging task. It is expected that more methods using spatial-temporal correlation will be proposed to solve this problem in the future.
Boosting effect of DaCM for ReID methods. Since the edge-level utilization in our method can be seen as one kind of reranking technologies, we embedded it to several visual ReID methods (PCB [30], SBS [10], TransReID [11]) to show the boosting performance. Methodologies for comparison can be categorized into several different groups, including several classical methods including LOMO+XQDA [20], handcrafted approach BoW+kissme [38], explicit deep learning methods (PAN [40], SVDNet [29] and HA-CNN), attribute-centric techniques APR [21], mask-guided strategies Human Parsing [18], part-based approaches (MultiScale [3] and PSE+ECN [28]), pose-oriented techniques PCB [30], and CLIP-ReID [19]. The results for comparison are shown in Table 2, underscore the noteworthy performance of our approach. Without bells and whistles, our method outperforms all of the existing methods on the DukeMTMC-reID dataset. In addition, the robustness of our methodology is further highlighted when employing the same visual stream method. For instance, integrated with SBS [10], our approach outperforms st-ReID [33], elevating the rank-1 accuracy from 93.7% to 96.7%, and boosting mAP from 82.8% to 89.8%.

### v2

@@ 4.2 Objective Function
In this part, we describe how to train the DaCM network. For the training set $\mathcal{D}^{\mathrm{train}}$, we randomly select two samples ($q$ and $d$) with the same label ID but different camera IDs, and their spatial-temporal information is denoted $\{l^{q},t^{q},c^{q}\}$ and $\{l^{d},t^{d},c^{d}\}$, respectively. According to the time difference $t^{d}-t^{q}$ and camera id $c^{q}$, we obtain the output $\mathbf{y}$ via the proposed network, and then we optimize the network via the expectation:
$\arg\max_{\theta}\mathbb{E}_{(q,d)\sim\mathcal{D}^{\mathrm{train}}}[P(y=c^{d}|t^{q},t^{d},c^{q};\theta)],$ (6)
where $\theta$ is the parameters of the network. In implementation, this expectation can be easily converted to a classification problem, and we adopt the cross-entropy loss function to generate the gradients:
$\ell=-\log\frac{\exp(\mathbf{y}_{c^{d}})}{\sum_{i}^{C}\exp(\mathbf{y}_{i})},$ (7)
where $y$ is the output of the DaCM network. The classification solves such a problem: determining the camera at which an object will appear again after a duration of $t^{d}-t^{q}$, starting from camera $c^{q}$. the network will acquire knowledge about the topology of devices deployed in the system and some characteristics of the target’s movement, thereby facilitating efficient inference.
@@ 4.3 Inference
As discussed in Sec. 3, to achieve efficient inference, the system should promptly return the desired image to the cloud server by generating appropriate $\mathbf{s}$ and $b$. In this section, we elaborate on how to accomplish this goal by using the DaCM network.
@@
$b^{i}=\frac{z^{i}}{\sum_{j}z^{j}}*B,~{}z^{i}=\phi\left(\frac{\mathbf{y}}{\gamma_{0}}\right)_{i}*\left(\frac{\exp(|\mathcal{G}^{i}|)}{\gamma_{1}\sum_{j}\exp(|\mathcal{G}^{j}|)}\right),$ (9)
where $\phi$ is the $\operatorname{softmax}$ function. Note that for cloud-level inference, $t^{d}$ should be provided by the user.
Edge-level inference. For edge-level inference, the focus is on re-ranking the index of gallery images in the uploading sequence by generating proper $\mathbf{s}^{i}$. Denoted the query image with its spatial-temporal information as $\{\mathcal{I}^{q},c^{q},t^{q}\}$, one gallery image at the $c^{d}$-th edge device as $\{\mathcal{I}^{d},c^{d},t^{d}\}$, we send $\{c^{q},t^{q},t^{d}\}$ into DaCM and obtain the output $\mathbf{y}$. If $\mathbf{y}_{c^{q}}$ is small, it implies that $\mathcal{I}^{q}$ and $\mathcal{I}^{d}$ do not match in spatial-temporal correlation, resulting in a minimal likelihood of having the same ID as the query image. Applying this operation to all gallery images on the $i$-th edge device, we obtain spatial-temporal similarity. Next, we delve into how to combine such spatial-temporal similarity with visual similarity.
Given a reliable visual similarity, it is difficult to build a reliable joint metric because the spatial-temporal similarity is unreliable and it is hard to assign appropriate weighting factors for these two types of metrics. Inspired by the joint metric proposed in (Wang et al. 2019), we adopt a smoothing operator to alleviate unreliable probability estimation. Denoted the spatial-temporal similarity as $\mathbf{o}\in\mathbb{R}^{N_{i}}$ and the visual similarity as $\mathbf{v}\in\mathbb{R}^{N_{i}}$ (assume it is produced by cosine distance function, and large value in $\mathbf{v}$ means the two features are similar), where $N_{i}$ is the number of gallery images in the $i$-th device, the joint similarity is computed as:
$s_{k}^{i}=-\frac{1}{1+\alpha\exp\left(\phi(-\frac{\mathbf{o}}{\beta})_{k}\right)}\frac{1}{1+\exp(\mathbf{v}_{k}-1)},$ (10)
where $\alpha$ and $\beta$ are two hyper-parameters to balance these two similarities, and the gallery images are re-ranked according to $\mathbf{s}^{i}$.
Time-constrained ReID. For a ReID system, users sometimes wish to search for targets near a specific time, a task challenging to accomplish solely based on visual features. Sorting only by time may introduce a large number of unrelated images. Therefore, the key to achieving tcReID lies in how to effectively combine time information with visual information. We observe that Eq. (10) offers a natural way to fulfill such a task. However, Eq. (10) does not satisfy the requirement because it does not introduce $t^{d}$ into the similarity calculation. Therefore, we propose a new formulation to meet the requirements of tcReID task.
@@
$\hat{s}_{k}^{i}=\frac{s_{k}^{i}}{1+\exp\left(\cos(\mathbf{B}_{k},\mathbf{y})-1\right)},$ (11)
where $\cos$ denotes the cosine distance function. Finally, we sort the gallery images according to $\hat{\mathbf{s}}^{i}$ and return them to the cloud server in batches. In addition, we find that the solely using the outputs of DaCM may lead to outlier problems. Therefore, in order to ensure the stability, we combine the output of DaCM with the spatial-temporal correlation obtained by the frequency statistics method (Wang et al. 2019).
@@ 5.1 Experimental Settings
Datasets. We mainly evaluate our proposed framework and method on the DukeMTMC-reID (Zheng, Zheng, and Yang 2017) and Market-1501 (Zheng et al. 2015) datasets, since they are annotated with high-quality timestamp.
Compared methods. We compare our method with several inference strategies to show its performance, including:
  • Pattern-C denotes the conventional centralized inference strategy, which collects all images captured by edge devices and conducts similarities calculations in the cloud server. We use it as the baseline to show the boosting effectiveness of different inference strategies.
  • Pattern-CE denotes a simple cloud-edge collaborative inference strategy: the amount of transmission is evenly distributed to each edge device and each edge device assigns the upload sequence according to the distance between the query image and gallery images. By comparing with this strategy, we can see the improvements brought by the proposed DaCM network.
Besides, we design a DaCM network to boost the efficiency of the ReID system, and we also replace the DaCM in the system with stReID (Wang et al. 2019), which uses frequency statistics to model spatial-temporal associations, and it interpolates the statistical spatial-temporal distribution into the similarity calculation process for person ReID.
Hyper-parameters: To train the DaCM network, we employ Adam (Kingma and Ba 2015) as the optimizer. The initial learning rate is set to 0.01 and is reduced by 10 for every 30 epochs. $\gamma_{0}$ and $\gamma_{1}$ are both set to 0.01 as the default. $\alpha$ and $\beta$ are bot set to 0.1. $\lambda$ in Eq. (2) is set to 10,000 as the default. $B$ is set to $3*C$, i.e., each edge device can upload an average of three images at a time.
@@ 5.2 Evaluation Protocols
We propose several novel protocols to show the performance of the cloud-edge collaborative inference. Let us initially provide a definition of the desired image, as the proposed protocols hinge upon this conceptual foundation. A desired image is a particular sample among the gallery images, sharing the same identity as the query image and possessing a timestamp in proximity to a given target time.
  • mean Transmission Number(mTN): it is a protocol used to present the efficiency of the method. For each pair of one query image and one gallery image with same identity ID, there exists one corresponding TN as shown in Eq.(1), and we average the TN of all pairs as mTN.
  • precise Rank@K (PR-K): it is calculated by checking whether top-k gallery images contain the desired image that has the same ID with the query image and is closest to the target time, so pR-K is a stricter protocol than R-K.
  • mean precise Rank (mpR): For the $i$-th query data, when pR-$k_{i}$ is successful but pR-($k_{i}$-1) is not successful, we record its precise Rank as $k_{i}$, and we average the $k_{i}$ of all query data as mpR.
5.3 Experimental Results
Boosting effect of DaCM for ReID methods. Since the edge-level inference in our method can be seen as one kind of re-ranking technologies, we embedded it to several visual ReID methods (PCB (Sun et al. 2018), SBS (He et al. 2023), TransReID (He et al. 2021)) to show the boosting performance. Methodologies for comparison can be categorized into several different groups, including several classical methods such as LOMO+XQDA (Liao et al. 2015) and handcrafted approach BoW+kissme (Zheng et al. 2015), explicit deep learning methods including PAN (Zheng, Zheng, and Yang 2019), SVDNet (Sun et al. 2017) and HA-CNN (Li, Zhu, and Gong 2018), attribute-centric techniques including APR (Lin et al. 2019), mask-guided strategies including Human Parsing (Kalayeh et al. 2018), part-based approaches like PSE+ECN (Sarfraz et al. 2018), pose-oriented techniques like PCB (Sun et al. 2018), and a recent work CLIP-ReID (Li, Sun, and Li 2023).
The results evaluated on DukeMTMC-reID dataset for comparison are shown in Table 1. Without bells and whistles, our method outperforms all existing methods on the DukeMTMC-reID dataset. In addition, the robustness of our methodology is further highlighted when employing the same visual stream method. For instance, integrated with SBS (He et al. 2023), our approach outperforms st-ReID (Wang et al. 2019), elevating the rank-1 accuracy from 95.4% to 96.7%, and boosting mAP from 83.0% to 89.8%. Besides, the results evaluated on Market1501 dataset for comparison are shown in Table 2, and our method still gain obvious improvements than the baselines in term of the R-1.
Table 1: Boosting Effect on DukeMTMC-reID dataset.
@@ 5.3 Experimental Results
BoW+kissme (Zheng et al. 2015) 25.1 12.2
LOMO+XQDA (Liao et al. 2015) 30.8 17.0
PAN (Zheng, Zheng, and Yang 2019) 71.6 51.5
SVDNet (Sun et al. 2017) 76.7 56.8
HA-CNN (Li, Zhu, and Gong 2018) 80.5 63.8
APR (Lin et al. 2019) 70.7 51.9
Human Parsing (Kalayeh et al. 2018) 84.4 71.0
PSE+ECN (Sarfraz et al. 2018) 85.2 79.8
CLIP-ReID (Li, Sun, and Li 2023) 90.0 80.7
PCB (Sun et al. 2018) 82.3 70.7
PCB + stReID (Wang et al. 2019) 94.3 84.0
PCB + InSTD (Ren et al. 2021) 92.7 86.1
PCB + Ours (OE) 96.2 89.5
SBS (He et al. 2023) 90.8 79.9
SBS + stReID 95.4 83.0
SBS + Ours (OE) 96.7 89.8
TranReID (He et al. 2021) 90.8 81.8
TranReID + stReID 96.2 88.6
TranReID + Ours (OE) 96.8 91.0
Table 2: Boosting Effect on Market-1501 dataset.
@@
BoW+kissme (Zheng et al. 2015) 44.4 20.8
PAN (Zheng, Zheng, and Yang 2019) 82.8 63.4
SVDNet (Sun et al. 2017) 82.3 62.1
HA-CNN (Li, Zhu, and Gong 2018) 91.2 75.7
APR (Lin et al. 2019) 84.3 64.7
Human Parsing (Kalayeh et al. 2018) 93.9 -
PSE+ECN (Sarfraz et al. 2018) 90.3 84.0
CLIP-ReID (Li, Sun, and Li 2023) 95.7 89.8
SBS (He et al. 2023) 95.8 89.0
SBS + stReID 96.1 86.8
SBS + Ours(OE) 96.4 88.2
TranReID (He et al. 2021) 95.2 89.0
TranReID + stReID 96.6 89.8
TranReID + Ours(OE) 96.9 89.0
Table 3: Performance of Different Inference Strategies Evaluated on DukeMTMC-reid Dataset.
Methods C CE OC OC+OE
@@
Table 4: Performance of Different Inference Strategies Evaluated on DukeMTMC-reid Dataset.
protocol C Ours(Lin) Ours(Sam) Ours
pR-1$\uparrow$ 0.94 20.38 23.37 44.74
@@
Efficiency of the proposed ReID system. We compare the proposed approach with the strategies introduced in Sec. 5.1, and the results are shown in Table 3 and Table 4, where C and CE denotes the Pattern-C and Pattern-CE, respectively. OC and OE denotes using cloud-level inference and edge-level inference. By analyzing the mTN values of different strategies, we can see that Pattern-C obtains a huge number of mTN since it requires uploading all images to the cloud server, and a simple cloud-edge collaborative framework (Pattern-CE) reduces mTN to 9.56, which saves much network traffic. Meanwhile, the results in the table also show that using DaCM network alone in the cloud server or in the edge devices can reduce the mTN to a certain extent (from 9.56 to 5.39 and 4.43, respectively), and the combination of them can lead to an optimal result. As for the protocols of pR-1 and mpR, most of the previous methods do not take it into consideration, and their methods only produce meaningless output. As shown in Table 4, if we use the visual similarity, it only achieves 0.94 pR-1. However, we can learn that our method can achieve 44.74 pR-1, which is still low but it demonstrates that the proposed approach can be applied to the challenging task. It is expected that more methods using spatial-temporal correlation will be proposed to solve this problem in the future.
Figure 5: Effects of various $\gamma_{0}$ (left) and $\gamma_{1}$ (right) for mTN.